a production process produces 4% defective parts. a sample of six parts from the production process is selected. what is the probability that the sample contains more than one defective parts
Answer #1

Probability of a part being defective (p)= 4% = 0.04

Probability of a part not being defective (q) = 1-0.04 = 0.96

Sample size = 6

This is a case of binomial distribution.

Probability that the sample contains more than one defective parts=

1 - ( Probability of 0 defective parts + Probability of 1 defective part)

= 1 - (6C0 * (0.04)0 * (0.96)6 + 6C1 * (0.04)1 * (0.94)5 )

= 1 - 0.9784

= 0.02155